OpenStudy (anonymous):

What value is added to both sides of the equation x2 – 8x = –10 in order to solve by completing the square? a. –8 b. 16 c. –4 d. 4 I think its a, am i right?

OpenStudy (mertsj):

One half of 8 is 4 and 4 squared is 16 so add 16 to both sides.

OpenStudy (anonymous):

wait what

OpenStudy (anonymous):

what mertsj said, with exactly that reasoning

OpenStudy (anonymous):

forget the number on the right hand side, it is irrelevant

OpenStudy (anonymous):

you have \[x^2-\color{red}{8}x\] \[\frac{8}{2}=4, 4^2=16\] add 16 to both sides get \[(x-4)^2=-10+16=6\]

OpenStudy (anonymous):

you should think it in english exactly as mertsj wrote it half of 8 is 4, 4 squared is 16

OpenStudy (anonymous):

oh i see, i guess i understand better with visual. Thanks a ton!

OpenStudy (anonymous):

\[x^2+12x=-5\] half of 12 is 6, 6 squared is 36 \[(x+6)^2=-5+36\] etc